# On energy bounds in asymptotically locally AdS spacetimes

### v1

@@ Abstract
This work considers positive energy theorems in asymptotically, locally AdS spacetimes. Particular attention is given to spacetimes where conformal infinity has compact, Einstein cross-sections admitting Killing or parallel spinors; a positive energy theorem is derived for such spacetimes in terms of geometric data intrinsic to the cross-section. This is followed by the first complete proofs of the BPS inequalities in (the bosonic sectors of) 4D and 5D minimal, gauged supergravity, including with magnetic fields. The BPS inequalities are proven for asymptotically AdS spacetimes, but also generalised to the aforementioned class of asymptotically, locally AdS spacetimes.
@@ 1 Introduction
Given the deep connections between Witten’s method and supergravity [16], another natural extension is to try prove BPS inequalities for (the bosonic sectors of) supergravity theories. This was realised very soon after Witten’s original work to prove global mass-charge inequalities in asymptotically flat spacetimes in four and five dimensions [17, 18]. While some results already exist along these lines [19, 20, 21] in the context of asymptotically AdS spacetimes - i.e. in gauged supergravity theories - the magnetic field is set to zero in [19] and a non-gauge-covariant connection is used in [20, 21], thereby leading to some unnatural assumptions and different results to the present work when incorporating magnetic fields [^2] This is discussed further in section 5.. This paper aims to build upon the literature by providing a more complete treatment of magnetic fields in the study of classical energy-charge inequalities with negative cosmological constant. This includes BPS inequalities with the standard asymptotically (globally) AdS asymptotics, but also a study of how BPS inequalities may be changed in the asymptotically, locally AdS case.
The paper begins in section 2 by setting up the formalism of asymptotically, locally AdS spacetimes and applying Witten’s method to prove a very general positive energy theorem - theorem 2.9. Section 3 follows with an analysis of the boundary geometry and specialises theorem 2.9 to boundaries with the properties discussed above. Section 4 illustrates these results through various example boundary geometries - namely a squashed $S^{7}$, the torus, the round sphere and the lens space, $L(p,1)$. Finally, section 5 presents the BPS inequalities. One additional appendix justifies that a particular Dirac operator can be inverted, as required for Witten’s method. The main results are theorems 2.9, 10, 3.5, 3.6 and corollary 5.2.1.
@@ 2 Set-up and general positive energy theorem
This definition can be derived by background subtraction from “first principles” using the procedure described in [23, 14] to ensure a well-defined Hamiltonian formulation. In the examples studied later, only $\mathrm{AdS}_{5}$ and its variations will actually have a non-zero $\bar{f}_{(n-1)mn}$.
@@
There is also a more subtle issue with background Killing spinors; $\varepsilon_{k}$ may only be defined in an open neighbourhood of the “boundary” at infinity or equation 2.4 may only have a solution in such a region. This in itself is not a problem because equation 2.4 will only really be required in an open neighbourhood of infinity, say $\mkern 1.8mu\overline{\mkern-1.8muM\mkern-1.8mu}\mkern 1.8mu$, and $\varepsilon_{k}$ can be extended to a spinor on all of $\Sigma_{t}$ by multiplying it with a smooth function that’s one near infinity but falls to zero within $\mkern 1.8mu\overline{\mkern-1.8muM\mkern-1.8mu}\mkern 1.8mu$. The problem is that $(\mkern 1.8mu\overline{\mkern-1.8muM\mkern-1.8mu}\mkern 1.8mu,\bar{g})$ may admit multiple spin structures and the spin structure which admits a non-zero solution, $\varepsilon_{k}$, may not be compatible with the spin structure on $(M,g)$. This is exactly the issue behind the AdS soliton [27], which will be discussed again in sections 4.2 and 5. But, in short, the proofs here will only work when a spin structure admitting a non-zero $\varepsilon_{k}$ on $\mkern 1.8mu\overline{\mkern-1.8muM\mkern-1.8mu}\mkern 1.8mu$ is compatible with a spin structure on $M$.
@@
Having established these preliminaries, it’s now possible to state the Lichnerowicz identity associated to the connections in the definition above. Note that a variant of the Lichnerowicz identity always underpins any Witten-style positive energy theorem.
@@
[…]
@@
[…]
@@
Each term in equation 2.34 can be calculated using equation 2.14. In particular, one finds
$g(e_{M},[\partial_{r},e_{N}])$ $=-\eta_{MN}+\bar{e}_{M}^{(\bar{f})m}\partial_{r}\big{(}\bar{e}_{N}^{(\bar{f})q}\big{)}\bar{f}_{mq}$
$\,\,\,\,\,\,\,+\frac{n-1}{2}\mathrm{e}^{-(n-1)r}\bar{e}_{M}^{(\bar{f})m}\bar{e}_{N}^{(\bar{f})n}(f_{(n-1)mn}-\bar{f}_{(n-1)mn})+O(\mathrm{e}^{-nr}),$ (2.35)
$\bar{g}(\bar{e}_{M},[\partial_{r},\bar{e}_{N}])$ $=-\eta_{MN}+\bar{e}_{M}^{(\bar{f})m}\partial_{r}\big{(}\bar{e}_{N}^{(\bar{f})n}\big{)}\bar{f}_{mn}$ (2.36)
and therefore
@@
up to $O(\mathrm{e}^{-nr})$. Substituting this back into equation 2.34 and simplifying yields
$-\frac{1}{2}(\omega_{1MA}-\mkern 1.8mu\overline{\mkern-1.8mu\omega\mkern-1.8mu}\mkern 1.8mu_{1MA})\varepsilon_{k}^{\dagger}\gamma^{A}\gamma^{M}\varepsilon_{k}$ $=\frac{n-1}{4}\mathrm{e}^{-(n-1)r}\bar{e}_{M}^{(\bar{f})m}\bar{e}_{0}^{(\bar{f})n}(f_{(n-1)mn}-\bar{f}_{(n-1)mn})\bar{\varepsilon}_{k}\gamma^{M}\varepsilon_{k}$
@@
to leading order. The $\mathrm{e}^{-(n-1)r}$ factor and $\varepsilon_{k}=O(\mathrm{e}^{r/2})$ mean everything else can be kept to $O(1)$; anything lower order will go to zero in equation 2.30. Thereofore ultimately,
@@ 3 Energy bounds and boundary geometry
[…]
@@
From [31], the round sphere is the only compact, Riemannian manifold admitting non-trivial solutions to $\hat{D}_{A}^{(h)}\hat{D}_{B}^{(h)}\hat{s}^{(\pm)}=-\delta_{AB}\hat{s}^{(\pm)}$. As that case has been explicitly excluded in this theorem, it must be that $\hat{s}^{(\pm)}=0$ and consequently $\hat{\xi}^{(\pm)A}=0$.

### v2

@@ Abstract
This work considers positive energy theorems in asymptotically, locally AdS spacetimes. Particular attention is given to spacetimes where conformal infinity has compact, Einstein cross-sections admitting Killing or parallel spinors; a positive energy theorem is derived for such spacetimes in terms of geometric data intrinsic to the cross-section. This is followed by the first complete proofs of the BPS inequalities in (the bosonic sectors of) 4D and 5D minimal, gauged supergravity, including with magnetic fields, provided the Maxwell field is exact. The BPS inequalities are proven for asymptotically AdS spacetimes, but also generalised to the aforementioned class of asymptotically, locally AdS spacetimes.
@@ 1 Introduction
Given the deep connections between Witten’s method and supergravity [16], another natural extension is to try prove BPS inequalities for (the bosonic sectors of) supergravity theories. This was realised very soon after Witten’s original work to prove global mass-charge inequalities in asymptotically flat spacetimes in four and five dimensions [17, 18]. Some results already exist along these lines [19, 20, 21, 22] in the context of asymptotically AdS spacetimes - i.e. in gauged supergravity theories. However, the magnetic field is set to zero in [19] and a non-gauge-covariant connection is used in [20, 21], thereby leading to some unnatural assumptions and different results to the present work when incorporating magnetic fields. [22] has the closest results to the present work, but still has some dependence on [20] in its reasoning [^2] These issues are all discussed further in section 5.. This paper aims to build upon the literature by providing a more complete treatment of magnetic fields in the study of classical energy-charge inequalities with negative cosmological constant. This includes BPS inequalities with the standard asymptotically (globally) AdS asymptotics, but also a study of how BPS inequalities may be changed in the asymptotically, locally AdS case.
The paper begins in section 2 by setting up the formalism of asymptotically, locally AdS spacetimes and applying Witten’s method to prove a very general positive energy theorem - theorem 2.9. Section 3 follows with an analysis of the boundary geometry and specialises theorem 2.9 to boundaries with the properties discussed above. Section 4 illustrates these results through various example boundary geometries - namely a squashed $S^{7}$, the torus, the round sphere and the lens space, $L(p,1)$. Finally, section 5 presents the BPS inequalities. One additional appendix justifies that a particular Dirac operator can be inverted, as required for Witten’s method. The main results are theorems 2.9, 11, 3.5, 3.6 and corollary 5.2.1.
@@ 1.1 Conventions
$(M,g)$ is a smooth, $n$-dimensional spacetime with mostly pluses metric signature. $a,b,\cdots$, $I,J,\cdots$, $M,N,\cdots$ and $A,B,\cdots$ denote vielbein indices running $\{0,1,\cdots,n-1\}$, $\{1,2,\cdots,n-1\}$, $\{0,2,3,\cdots,n-1\}$ and $\{2,3,\cdots,n-1\}$ respectively [^3] Although in most cases, equations with $a,b,\cdots$ will remain valid if these were abstract indices. while $\mu,\nu,\cdots$, $i,j,\cdots$, $m,n,\cdots$ and $\alpha,\beta,\cdots$ denote coordinate indices running $\{0,1,\cdots,n-1\}$, $\{1,2,\cdots,n-1\}$, $\{0,2,3,\cdots,n-1\}$ and $\{2,3,\cdots,n-1\}$ respectively. It’s assumed $n\geq 4$ throughout. The gamma matrices are chosen so that $\gamma^{a}\gamma^{b}+\gamma^{b}\gamma^{a}=-2g^{ab}I$, $(\gamma^{0})^{\dagger}=\gamma^{0}$ and $(\gamma^{I})^{\dagger}=-\gamma^{I}$. $\gamma^{a_{1}\cdots a_{k}}$ denotes the antisymmetrisation, $\gamma^{[a_{1}}\cdots\gamma^{a_{k}]}$. Only Dirac spinors will be used in this work and the Dirac conjugate of a Dirac spinor, $\Psi$, is $\overline{\Psi}=\Psi^{\dagger}\gamma^{0}$. From section 3 onwards, it will be necessary to consider both Dirac spinors on an (n2)(n-2)-dimensional surface and Dirac spinors on the full spacetime. Thus, the spinors on the (n2)(n-2)-dimensional surface exist in a vector space which has half the dimension of the spacetime spinors. In these cases, the spacetime spinor space can be viewed as a direct sum of the (n2)(n-2)-dimensional surface’s spinor space with itself. Furthermore, the gamma matrices for the spacetime will be chosen as
@@ 2 Set-up and general positive energy theorem
This definition can be derived by background subtraction from “first principles” using the procedure described in [24, 14] to ensure a well-defined Hamiltonian formulation. In the examples studied later, only $\mathrm{AdS}_{5}$ and its variations will actually have a non-zero $\bar{f}_{(n-1)mn}$. The full geometric set-up described here is illustrated in figure 1.
Figure 1: This is a Penrose diagram for defining asymptotically, locally AdS spacetimes and energy within them. Conformal infinity, $\mathcal{I}$, is $\{r=\infty\}$, where $r$ is the Fefferman-Graham coordinate. $\mathcal{I}$ is topologically $\mathbb{R}\times S$ with metric $f_{(0)mn}$. $t$ is a coordinate along the $\mathbb{R}$ direction, $\Sigma_{t}$ is a spacelike hypersurface which has constant $t$ near $\mathcal{I}$ and $n^{a}$ is the future-directed, timelike, unit normal to $\Sigma_{t}$. Energy is measured at $\Sigma_{t,\infty}=\mathcal{I}\cap\Sigma_{t}$.
@@
Note that $\varepsilon_{k}$ may only be defined in an open neighbourhood of the “boundary” at infinity or equation 2.4 may only have a solution in such a region. This is not a problem because equation 2.4 will only really be required in an open neighbourhood of infinity, say $\mkern 1.8mu\overline{\mkern-1.8muM\mkern-1.8mu}\mkern 1.8mu$, and $\varepsilon_{k}$ can be extended to a spinor on all of $\Sigma_{t}$ by multiplying it with a smooth function that’s one near infinity but falls to zero within $\mkern 1.8mu\overline{\mkern-1.8muM\mkern-1.8mu}\mkern 1.8mu$.
@@
Having established these preliminaries, it’s now possible to state the Lichnerowicz identity associated to the connections in the definition above [^9] A similar Lichnerowicz identity is also given in [28].. Note that a variant of the Lichnerowicz identity always underpins any Witten-style positive energy theorem.
@@
$=\delta_{M0}\hat{f}_{(0)}^{mn}(f_{(n-1)mn}-\bar{f}_{(n-1)mn})+\delta^{{{A}\mathchoice{\makebox[8.3764pt][c]{$\displaystyle$}}{\makebox[8.3764pt][c]{$\textstyle$}}{\makebox[5.2879pt][c]{$\scriptstyle$}}{\makebox[3.77707pt][c]{$\scriptscriptstyle$}}}}_{{\mathchoice{\makebox[6.01392pt][c]{$\displaystyle$}}{\makebox[6.01392pt][c]{$\textstyle$}}{\makebox[3.675pt][c]{$\scriptstyle$}}{\makebox[2.625pt][c]{$\scriptscriptstyle$}}{M}}}e_{A}^{(f_{(0)})m}n_{(0)}^{n}(f_{(n-1)mn}-\bar{f}_{(n-1)mn}).$ (2.24)
@@
$D_{A}\varepsilon_{k}$ $=(e^{{\mathchoice{\makebox[6.01392pt][c]{$\displaystyle$}}{\makebox[6.01392pt][c]{$\textstyle$}}{\makebox[3.675pt][c]{$\scriptstyle$}}{\makebox[2.625pt][c]{$\scriptscriptstyle$}}{m}}}_{{{A}\mathchoice{\makebox[7.09612pt][c]{$\displaystyle$}}{\makebox[7.09612pt][c]{$\textstyle$}}{\makebox[4.30226pt][c]{$\scriptstyle$}}{\makebox[3.07304pt][c]{$\scriptscriptstyle$}}}}-\bar{e}^{{\mathchoice{\makebox[6.01392pt][c]{$\displaystyle$}}{\makebox[6.01392pt][c]{$\textstyle$}}{\makebox[3.675pt][c]{$\scriptstyle$}}{\makebox[2.625pt][c]{$\scriptscriptstyle$}}{m}}}_{{{A}\mathchoice{\makebox[7.09612pt][c]{$\displaystyle$}}{\makebox[7.09612pt][c]{$\textstyle$}}{\makebox[4.30226pt][c]{$\scriptstyle$}}{\makebox[3.07304pt][c]{$\scriptscriptstyle$}}}})\partial_{m}\varepsilon_{k}-\frac{1}{4}(\omega_{abA}-\mkern 1.8mu\overline{\mkern-1.8mu\omega\mkern-1.8mu}\mkern 1.8mu_{abA})\gamma^{ab}\varepsilon_{k}-\frac{\mathrm{i}}{2}\gamma_{A}\varepsilon_{k}.$ (2.31)
@@
Each term in equation 2.34 can be calculated using equation 2.14. In particular,
$g(e_{M},[\partial_{r},e_{N}])$
$=g\bigg(\mathrm{e}^{-r}\bar{e}_{M}^{(\bar{f})m}\partial_{m}-\frac{1}{2}\mathrm{e}^{-nr}\bar{e}_{M}^{(\bar{f})m}(f_{(n-1)mp}-\bar{f}_{(n-1)mp})\bar{f}^{pn}_{(0)}\partial_{n}+O(\mathrm{e}^{-(n+1)r}),$
$\,\,\,\,\,\,\,-\mathrm{e}^{-r}\bar{e}_{N}^{(\bar{f})q}\partial_{q}+\mathrm{e}^{-r}\partial_{r}\big(\bar{e}_{N}^{(\bar{f})q}\big)\partial_{q}+\frac{n}{2}\mathrm{e}^{-nr}\bar{e}_{N}^{(\bar{f})q}(f_{(n-1)qr}-\bar{f}_{(n-1)qr})\bar{f}^{rs}_{(0)}\partial_{s}+O(\mathrm{e}^{-(n+1)r})\bigg)$ (2.35)
$=\bar{e}_{M}^{(\bar{f})m}\partial_{r}\big(\bar{e}_{N}^{(\bar{f})q}\big)\Big(\bar{f}_{mq}+O(\mathrm{e}^{-(n-1)r})\Big)$
$\,\,\,\,\,\,\,-\bar{e}_{M}^{(\bar{f})m}\bar{e}_{N}^{(\bar{f})q}\Big(\bar{f}_{mq}+\mathrm{e}^{-(n-1)r}(f_{(n-1)mq}-\bar{f}_{(n-1)mq})-\frac{1}{2}\mathrm{e}^{-(n-1)r}(f_{(n-1)mp}-\bar{f}_{(n-1)mp})\bar{f}^{pn}_{(0)}\bar{f}_{(0)nq}$
$\,\,\,\,\,\,\,-\frac{n}{2}\mathrm{e}^{-(n-1)r}(f_{(n-1)qr}-\bar{f}_{(n-1)qr})\bar{f}^{rs}_{(0)}\bar{f}_{(0)ms}+O(\mathrm{e}^{-nr})\Big)$ (2.36)
$=\bar{e}_{M}^{(\bar{f})m}\partial_{r}\big(\bar{e}_{N}^{(\bar{f})q}\big)\bar{f}_{mq}+O(\mathrm{e}^{-nr})-\eta_{MN}$
$\,\,\,\,\,\,\,+\frac{n-1}{2}\mathrm{e}^{-(n-1)r}\bar{e}_{M}^{(\bar{f})m}\bar{e}_{N}^{(\bar{f})n}(f_{(n-1)mn}-\bar{f}_{(n-1)mn})+O(\mathrm{e}^{-nr})$ (2.37)
and likewise
$\bar{g}(\bar{e}_{M},[\partial_{r},\bar{e}_{N}])$ $=\bar{g}\left(\mathrm{e}^{-r}\bar{e}_{M}^{(\bar{f})m}\partial_{m},-\mathrm{e}^{-r}\bar{e}_{N}^{(\bar{f})n}\partial_{n}+\mathrm{e}^{-r}\partial_{r}\big(\bar{e}_{N}^{(\bar{f})n}\partial_{n}\big)\right)$ (2.38)
$=-\eta_{MN}+\bar{e}_{M}^{(\bar{f})m}\partial_{r}\big(\bar{e}_{N}^{(\bar{f})n}\big)\bar{f}_{mn}.$ (2.39)
Putting them together yields
@@
up to $O(\mathrm{e}^{-nr})$. Substituting this back into equation 2.34 then implies
$-\frac{1}{2}(\omega_{1MA}-\mkern 1.8mu\overline{\mkern-1.8mu\omega\mkern-1.8mu}\mkern 1.8mu_{1MA})\varepsilon_{k}^{\dagger}\gamma^{A}\gamma^{M}\varepsilon_{k}$ $=\delta^{AB}\frac{n-1}{4}\mathrm{e}^{-(n-1)r}\bar{e}_{A}^{(\bar{f})m}\bar{e}_{B}^{(\bar{f})n}(f_{(n-1)mn}-\bar{f}_{(n-1)mn})\varepsilon_{k}^{\dagger}\varepsilon_{k}$
$\,\,\,\,\,\,\,+\frac{n-1}{4}\mathrm{e}^{-(n-1)r}\bar{e}_{A}^{(\bar{f})m}\bar{e}_{0}^{(\bar{f})n}(f_{(n-1)mn}-\bar{f}_{(n-1)mn})\bar{\varepsilon}_{k}\gamma^{A}\varepsilon_{k}$ (2.41)
$=\frac{n-1}{4}\mathrm{e}^{-(n-1)r}\bar{e}_{M}^{(\bar{f})m}\bar{e}_{0}^{(\bar{f})n}(f_{(n-1)mn}-\bar{f}_{(n-1)mn})\bar{\varepsilon}_{k}\gamma^{M}\varepsilon_{k}$
@@
to leading order. The $\mathrm{e}^{-(n-1)r}$ factor and $\varepsilon_{k}=O(\mathrm{e}^{r/2})$ mean everything else can be kept to $O(1)$; anything lower order will go to zero in equation 2.30. Therefore ultimately,
@@ 3 Energy bounds and boundary geometry
$e^{{\mathchoice{\makebox[4.33765pt][c]{$\displaystyle$}}{\makebox[4.33765pt][c]{$\textstyle$}}{\makebox[2.59009pt][c]{$\scriptstyle$}}{\makebox[1.85005pt][c]{$\scriptscriptstyle$}}{\mu}}}_{{{a}\mathchoice{\makebox[4.86232pt][c]{$\displaystyle$}}{\makebox[4.86232pt][c]{$\textstyle$}}{\makebox[2.95248pt][c]{$\scriptstyle$}}{\makebox[2.10892pt][c]{$\scriptscriptstyle$}}}}\partial_{\mu}\varepsilon_{k}-\frac{1}{4}\omega_{bca}\gamma^{bc}\varepsilon_{k}+\frac{\mathrm{i}}{2}\gamma_{a}\varepsilon_{k}=0.$ (3.3)
@@
By Obata’s theorem [32], the round sphere is the only compact, Riemannian manifold admitting non-trivial solutions to $\hat{D}_{A}^{(h)}\hat{D}_{B}^{(h)}\hat{s}^{(\pm)}=-\delta_{AB}\hat{s}^{(\pm)}$. As that case has been explicitly excluded in this theorem, it must be that $\hat{s}^{(\pm)}=0$ and consequently $\hat{\xi}^{(\pm)A}=0$.
@@
To better understand the full physical significance of the inequalities just derived, it’s worth considering the geometric invariance of $E$ and $Q_{\hat{k}}$. In particular, $\mathcal{I}$ is defined through a conformal compactification and choosing a different compactification could change $f_{(0)}$ or $f_{(n-1)}$ through a change to the Fefferman-Graham parameter, $r$. It will be easiest to see this by swapping $r$ for $z=\mathrm{e}^{-r}$. Then,
$g=\frac{1}{z^{2}}\left(\mathrm{d}z\otimes\mathrm{d}z+(f_{(0)mn}+zf_{(1)mn}+\cdots)\mathrm{d}x^{m}\otimes\mathrm{d}x^{n}\right)$ (3.44)
and the conformal factor for compactification is $\Omega=z$. The most general transformations preserving the Fefferman-Graham form are known as Penrose-Brown-Henneaux (PBH) transformations and were derived in [33]. In particular, the new coordinates are defined by
$z=z^{\prime}\mathrm{e}^{-\sigma(x^{\prime})}\,\,\,\mathrm{and}\,\,\,x^{m}=x^{\prime m}+\xi^{m}(x^{\prime},z^{\prime})$ (3.45)
for an arbitrary function, $\sigma(x^{\prime})$. Then, $\xi^{m}$ is completely determined; namely it takes the form,
$\xi^{m}(x^{\prime},z^{\prime})$ $=\xi^{m}(x^{\prime},0)+\partial^{\prime}_{n}(\sigma)\int_{0}^{z^{\prime}}\zeta f^{mn}(x^{\prime},\zeta)\mathrm{d}\zeta.$ (3.46)
Practically, it suffices to consider $\sigma(x^{\prime})$ (and hence $\xi^{m}(x^{\prime},z^{\prime})$) to be infinitesimal. Then, the metric changes by
$\delta f$ $=\mathcal{L}_{\xi}f+2\sigma f-\sigma z\partial_{z}f.$ (3.47)
Most importantly for the present discussion, the boundary metric changes as
$\delta f_{(0)}$ $=\mathcal{L}_{\xi|_{z=0}}f_{(0)}+2\sigma f_{(0)}.$ (3.48)
Therefore, in general, the boundary metric can change in an arbitrary way. Furthermore, by expanding $\xi^{m}$ in powers of $z$, one can also find a complicated transformation law for $\delta f_{(n-1)}$ (separately for each value of $n$). In summary, the boundary data, $(f_{(0)},f_{(n-1)})$, could change in a very complicated way and therefore affect the expressions for $E$ and $Q_{\hat{k}}$.
However, this is an issue that affects even the standard asymptotically AdS spacetimes (i.e. with round sphere cross-section) and doesn’t seem to have been considered in any detail previously in the literature. Firstly, it’s unclear how physically meaningful this question is. For example, even an asymptotically flat end can be written in a poor choice of coordinates, which then leads to strange results for ADM mass [34]. Therefore, it would seem reasonable to insist on a particular conformal class representative for $f_{(0)}$. Indeed, this is what is done in section 5.3 of [35] when studying the first law of black hole mechanics in this context.
A more tractable, and physically well-motivated, problem regarding geometric invariance is to study the question considered in [11] for the energy in asymptotically AdS spacetimes [^13] The analogous problems for for asymptotically hyperbolic Riemannian manifolds and the ADM energy are studied in [9] and [36, 37] respectively.. Specifically, the task is to quantify how $E$ and $Q_{\hat{k}}$ change under asymptotic symmetries preserving both $\Sigma_{t}$ and the Fefferman-Graham gauge.
An asymptotic symmetry is determined by a vector field, $\xi$, such that $(\mathcal{L}_{\xi}g)|_{r=\infty}=0$. Meanwhile, following section 3 of [11], to preserve $\Sigma_{t}$, choose $\xi=\xi^{i}\partial_{i}=\xi^{1}\partial_{r}+\xi^{\alpha}\partial_{\alpha}$. Then,
$(\mathcal{L}_{\xi}g)_{\mu\nu}$ $=\xi^{1}\partial_{r}g_{\mu\nu}+g_{1\nu}\partial_{\mu}\xi^{1}+g_{\mu 1}\partial_{\nu}\xi^{1}+\xi^{\alpha}\partial_{\alpha}g_{\mu\nu}+g_{\alpha\nu}\partial_{\mu}\xi^{\alpha}+g_{\mu\alpha}\partial_{\nu}\xi^{\alpha}.$ (3.49)
The various components then reduce to
$(\mathcal{L}_{\xi}g)_{11}$ $=2\partial_{r}\xi^{1},$ (3.50)
$(\mathcal{L}_{\xi}g)_{1m}$ $=\partial_{m}\xi^{1}+\mathrm{e}^{2r}f_{m\alpha}\partial_{r}\xi^{\alpha}\,\,\,\mathrm{and}$ (3.51)
$(\mathcal{L}_{\xi}g)_{mn}$ $=\xi^{1}\partial_{r}(\mathrm{e}^{2r}f_{mn})+\mathrm{e}^{2r}(\xi^{\alpha}\partial_{\alpha}f_{mn}+f_{\alpha n}\partial_{m}\xi^{\alpha}+f_{m\alpha}\partial_{n}\xi^{\alpha}).$ (3.52)
$\xi$ should be a physical spacetime vector that extends to the boundary. Hence, in the spirit of the Fefferman-Graham expansion, let
$\xi=\xi_{(0)}+\mathrm{e}^{-r}\xi_{(1)}+\mathrm{e}^{-2r}\xi_{(2)}+\cdots$ (3.53)
for $\xi_{(k)}$ being $r$-independent. Then, for a spacetime that’s asymptotically AdS with cross-section, $(S,h)$, one finds
$(\mathcal{L}_{\xi}g)_{11}$ $=-2(\mathrm{e}^{-r}\xi_{(1)}+\cdots)$ (3.54)
$(\mathcal{L}_{\xi}g)_{1m}$ $=-\mathrm{e}^{r}f_{(0)m\alpha}\xi^{\alpha}_{(1)}+\partial_{m}(\xi^{1}_{(0)})-2f_{(0)m\alpha}\xi^{\alpha}_{(2)}+\cdots\,\,\,\mathrm{and}$ (3.55)
$(\mathcal{L}_{\xi}g)_{mn}$ $=\mathrm{e}^{2r}\left(2\xi^{1}_{(0)}f_{(0)mn}+\xi^{\alpha}_{(0)}\partial_{\alpha}f_{(0)mn}+f_{(0)\alpha n}\partial_{m}\xi^{\alpha}_{(0)}+f_{(0)m\alpha}\partial_{n}\xi^{\alpha}_{(0)}\right)+2\mathrm{e}^{r}\xi^{1}_{(1)}f_{(0)mn}$
$\,\,\,\,\,\,\,+\xi^{\alpha}_{(0)}\partial_{\alpha}f_{(2)mn}+f_{(2)\alpha n}\partial_{m}\xi^{\alpha}_{(0)}+f_{(2)m\alpha}\partial_{n}\xi^{\alpha}_{(0)}$
$\,\,\,\,\,\,\,+\xi^{\alpha}_{(2)}\partial_{\alpha}f_{(0)mn}+f_{(0)\alpha n}\partial_{m}\xi^{\alpha}_{(2)}+f_{(0)m\alpha}\partial_{n}\xi^{\alpha}_{(2)}+\cdots.$ (3.56)
Therefore, $(\mathcal{L}_{\xi}g)_{11}|_{r=\infty}$ holds automatically. Next, $(\mathcal{L}_{\xi}g)_{1m}|_{r=\infty}=0$ implies $f_{(0)m\alpha}\xi^{\alpha}_{(1)}=0$ and $\partial_{m}(\xi^{1}_{(0)})-2f_{(0)m\alpha}\xi^{\alpha}_{(2)}=0$. Since $f_{(0)mn}$ is invertible, the former condition says $\xi^{\alpha}_{(1)}=0$. Meanwhile, the latter condition says $\partial_{t}\xi^{1}_{(0)}=0$ and $2\xi_{(2)\alpha}=\partial_{\alpha}\xi^{1}_{(0)}$, where the index has been lowered using $h$. Finally, consider $(\mathcal{L}_{\xi}g)_{mn}|_{r=\infty}$. From the $\mathrm{e}^{r}$ component, it follows that $\xi^{1}_{(1)}=0$, meaning $\xi_{(1)}=0$ when combined with $\xi^{\alpha}_{(1)}=0$ from earlier. From the $\mathrm{e}^{2r}$ and $r$-independent components, setting $(m,n)=(\alpha,\beta)$ implies
$\xi^{1}_{(0)}h=-\frac{1}{2}\mathcal{L}_{\xi_{(0)}}h\,\,\,\mathrm{and}\,\,-\frac{c}{2}\mathcal{L}_{\xi_{(0)}}h+\mathcal{L}_{\xi_{(2)}}h=0,$ (3.57)
where the vector field argument of the Lie derivatives only considers the part tangential to $S$. Re-writing the Lie derivatives in terms of $\hat{D}^{(h)}$ and applying $2\xi_{(2)\alpha}=\partial_{\alpha}\xi^{1}_{(0)}$ from above, it then follows that
$\hat{D}^{(h)}_{\alpha}\hat{D}^{(h)}_{\beta}\xi^{1}_{(0)}=-c\xi^{1}_{(0)}h_{\alpha\beta}.$ (3.58)
Since $(S,h)$ is assumed to not be the round sphere [^14] The case of the round sphere will be considered later in section 4.3. in theorems 3.5 and 3.6, by the Obata theorem [32], it must be that $\xi^{1}_{(0)}=0$. Substituting back into equation 3.57 implies $\mathcal{L}_{\xi_{(0)}}h=0$, i.e. $\xi_{(0)}$ is a just a Killing vector of $h$.
Therefore $\delta f_{(0)\alpha\beta}=\delta h_{\alpha\beta}=0$ (by virtue of being an asymptotic symmetry) and $\mathcal{L}_{\xi_{(0)}}h=0$. Hence, $\sigma=0$ by equation 3.48. Being an asymptotic symmetry only determines the leading order behaviour of $\xi$. However, there is still the additional requirement of staying in Fefferman-Graham gauge. From equation 3.45 with $\sigma=0$, it follows that $\xi=\xi^{\alpha}_{(0)}\partial_{\alpha}$.
In summary, the only asymptotic symmetries in this context are coordinate transformations generated by the Killing vectors of $(S,h)$. These don’t affect $f_{(0)}$ by definition. Moreover, since there is no $z$ dependence, they also don’t affect the Taylor series split for $f_{mn}$; in particular $f_{(n-1)mn}$ just transforms as a tensor under coordinate transformations. Since the integrands of $E$ and $Q_{\hat{k}}$ are scalars, they remain invariant under these transformations.